OpenStudy (anonymous):

What is the smallest measure of the triangle if the interior angles are represented by 2x+10, 5x-30, and 3x+40?

OpenStudy (hba):

Do you know what is the sum of angles ?

OpenStudy (anonymous):

No, the answer choices are (a) 50 (b) 88 (c) 42 (d) 16

OpenStudy (hba):

I am not telling the answer Sum of interior angles =(n-2)*180 Here n=3

OpenStudy (hba):

Now solve.

OpenStudy (anonymous):

I still dont understand

OpenStudy (anonymous):

@Koda yeah I'm not getting where HBA is coming from either.

OpenStudy (anonymous):

yea, I'm really confused

OpenStudy (hba):

@BrittanyFrench Allow me to show.

OpenStudy (anonymous):

well okay then..

OpenStudy (hba):

Sum of interior angles =(3-2)*180 Sum of interior angles =180 Now 2x+10 + 5x-30 + 3x+40=180 Now solve for x.

OpenStudy (hba):

@Koda Now solve for x.

OpenStudy (anonymous):

@Koda , do you get that? I understnad now..

OpenStudy (anonymous):

thanks I get it now

OpenStudy (hba):

You are welcome :)

OpenStudy (anonymous):

Glad you underdstand!